[Brahms-l] Shift report 20050421 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Apr 21 2005 - 08:08:53 EDT
Shifter leader was kh
Shifters were 


Had beam the whole night.  Acquired the following runs
FS@861B, MRS 1050B
Run   trig2    trig3 
14159  135k     108k
FS@1219B, MRS 1050B
14162  125k      86k
14163   96k     129k
14167   81k     111k
14168   75k     101k
14175   16k      21k
14176   39k      53k

The exclusive trigger 5 quit working sometime yesterday.  This was 
eventually fixed by cycling the power on the sc vme crate.  After
that the spin scaler scripts did not work.  Flemming fixed that from home by rebooting that processor.  By the end of the shift, all appears to be working

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Apr 21 08:09:36 2005

This archive was generated by hypermail 2.1.8 : Thu Apr 21 2005 - 08:09:53 EDT